Alexander an artist, welded 3 pieces of bronze together to create a project. The second piece is one-half of the first piece. The third piece is 4 less that the first piece. The total length of the pieces is 86 inches. How long was each piece?
Answer by oberobic(2304)

F = length of first piece
S = length of second piece = .5F
T = length of third piece = F-4
F+S+T=86
Substituting what we know...
F + .5F + (F-4) = 86
Combining...
2.5F - 4 = 86
Adding 4 to both sides...
2.5F = 90
Dividing both sides by 2.5
F = 36
Substituting back into our formulas...
S = .5F = .5(36) = 18
T = F - 4 = 36 - 4 = 32
Check the sum...
36 + 18 + 32 (equals?) 86?
Yes it does.